# Fully heavy tetraquark states with diquark-antidiquark configuration

### v1

@@ III.1 Model parameters
$M=$ $2m_{q_{1}}+m_{q_{2}}+(\kappa_{q_{1}q_{1}})_{\bar{3}}[S(S+1)-\frac{3}{2}]$
$+$ $(\kappa_{q_{1}q_{2}})_{\bar{3}}[J(J+1)-S(S+1)-\frac{3}{4}],$ (14)
where $J$ is the $S$ represents the spin of two same-flavor quarks within a baryon.
By considering baryons composed of nnqnnq with quantum numbers $(S,J)$ equal to $(0,\frac{1}{2})$, $(1,\frac{1}{2})$, and $(1,\frac{3}{2})$, the values of ${(\kappa_{nn})}_{\bar{3}}$ and ${(\kappa_{nq})}_{\bar{3}}$ can be determined. Here $n=u,d$ and $q=s,c,b$. We can get a proportional relationship read ${(\kappa_{nn})}_{\bar{3}}/{(\kappa_{nq})}_{\bar{3}}=m_{q}/m_{n}$. Furthermore, the values of ${(\kappa_{bb})}_{\bar{3}}$, ${(\kappa_{cb})}_{\bar{3}}$ can be extracted from $\Omega_{bbb}$, $\Omega_{bbc}$ and $\Omega_{bbc}^{*}$. We can extract ${(\kappa_{cc}})_{\bar{3}}$ from $\Omega_{ccc}$, $\Omega_{ccs}$ and $\Omega_{ccs}^{*}$. All the coefficients are shown in Table 1.
At this juncture, it remains crucial to determine the value of $(\kappa_{QQ})_{6}$, along with establishing the corresponding $(\kappa_{Q\bar{Q}})_{T}$ parameter within the tetraquark configuration. From one-gluon-exchange, we obtain
$(\kappa_{QQ})_{6}={}$ $-\frac{1}{2}(\kappa_{QQ})_{\bar{3}},$ (15)
$(\kappa_{Q\bar{Q}})_{\bar{3}}={}$ $\frac{1}{4}(\kappa_{Q\bar{Q}})_{1},$
$(\kappa_{Q\bar{Q}})_{6}={}$ $\frac{5}{8}(\kappa_{Q\bar{Q}})_{1}.$
where $(\kappa_{Q\bar{Q}})_{\bar{3}}$ means the $\kappa_{Q\bar{Q}}$ for quark-antiquark in tetraquark $|(Q_{1}Q_{2})^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})^{3}\rangle$ and $(\kappa_{Q\bar{Q}})_{6}$ for $|(Q_{1}Q_{2})^{6}(\bar{Q}_{3}\bar{Q}_{4})^{\bar{6}}\rangle$.
The masses of charm and bottom quark we derived are $m_{c}=1556$MeV and $m_{b}=4755$MeV.
After computing the ground state masses, we can include the contributions from $H_{SL}$ and $H_{L}$ to calculate the masses of the $L=1$ excited states. By applying Eq. (1) to charmonium mesons with $S=1$ and $L=0,1$, specifically $J/\psi$, $\chi_{c0}$, $\chi_{c1}$, and $\chi_{c2}$, we can fit the coefficients in $H_{SL}$ and $H_{L}$, we get $A_{cc}=21.83$ MeV and $B_{cc}=426.06$ MeV. Similarly, using the bottomonium mesons $\Upsilon$, $\chi_{b0}$, $\chi_{b1}$, and $\chi_{b2}$, we obtain $A_{bb}=8.23$ MeV and $B_{bb}=438.72$ MeV. For tetraquark states containing both charm and bottom quarks, we use the average values of $A_{cc}$ and $A_{bb}$ (as well as $B_{cc}$ and $B_{bb}$) for our calculations.
Since the tetraquark states discussed in this work are structured as diquark–antidiquark systems, quark rearrangement is required for them to strongly decay into two mesons. The decay widths are calculated using Eq. (11) with $\mathcal{A}$ = 2600 MeV. The value of $\mathcal{A}$ can be obtained by fitting light scalar meson decays [57]. As higher-order decays (such as $P$-wave and $D$-wave decays) contribute less significantly to the total decay width, our analysis focuses primarily on $S$-wave decays. However, we also analyze and discuss $P$-wave decays in special cases.
@@
Figure 1: The mass spectra of the fully-heavy tetraquark states: (a) $cc][\bar{c}\bar{c}]$, (b) $bb][\bar{b}\bar{b}]$, and (c) $cb][\bar{c}\bar{b}]$, each with different quantum numbers. The dash-dotted lines represent the corresponding meson-meson thresholds. In Fig. (c), the “several channels” include many threshold lines; for clarity and visual neatness, those lines overlapping with the energy levels have been omitted.
@@
Table 2: The quantum numbers, masses, eigenvectors of the basis wave functions, decay channels of $cc][\bar{c}\bar{c}]$ and $bb][\bar{b}\bar{b}]$ configurations
@@
Table 3: The quantum numbers, masses, eigenvectors of the basis wave functions and decay channels of $cb][\bar{c}\bar{b}]$ configuration. For certain states, there are no available channels that both satisfy the conservation laws and have thresholds below the mass of the state.
@@
Table 4: The quantum numbers, masses, eigenvectors of the basis wave functions and decay channels of $cc][\bar{c}\bar{b}]$ configuration
@@
Table 5: The quantum numbers, masses, eigenvectors of the basis wave functions and decay channels of $bb][\bar{b}\bar{c}]$ configuration
@@
Table 6: The quantum numbers, masses, eigenvectors of the basis wave functions and decay channels of $cc][\bar{b}\bar{b}]$ configuration. For certain states, there are no available channels that both satisfy the conservation laws and have thresholds below the mass of the state.
@@
III.2 $cc][\bar{c}\bar{c}]$, $bb][\bar{b}\bar{b}]$and $cb][\bar{c}\bar{b}]$ systems
From the wave functions of $cc][\bar{c}\bar{c}]$, $bb][\bar{b}\bar{b}]$ and $cb][\bar{c}\bar{b}]$ configurations, it can be seen that these states possesses a well-defined charge conjugation parity ($C$-parity). For these states, conservation of $C$-parity must be taken into account during the decay process. Their spectra and decay channels were shown in Fig. 1 and Tables 2-3. The final-state mesons we choose for these decays are in the $1S$ and $1P$ states. To date, only the spin-0 states of the $B_{c}(1S)$ and $B_{c}(2S)$ mesons have been observed experimentally. Therefore, for the remaining $c\bar{b}$ mesons in the $1S$ and $1P$ states, we use the masses predicted by lattice QCD calculations [58]. For simplicity, we denote the $c\bar{b}$ mesons with quantum numbers $J^{PC}=0^{+}$ and $1^{+}$ as $B_{c0}$ and $B_{c1}$, respectively.
@@ III.2.1 $cc][\bar{c}\bar{c}]$ system
The $S$-wave fully charmed tetraquark states possess quantum numbers $J^{PC}=0^{++}$, $1^{+-}$, and $2^{++}$. There are two states with $0^{++}$, with masses of 6098.45 MeV and 6314.21 MeV, respectively. The lower-mass state, at 6098.45 MeV, lies above both the $\eta_{c}\eta_{c}$ and $\eta_{c}J/\psi$ thresholds. However, due to conservation laws, it can only decay naturally into the $\eta_{c}\eta_{c}$ channel, while decay into $\eta_{c}J/\psi$ is forbidden. The higher-mass $0^{++}$ state, at 6314.21 MeV, lies above both $\eta_{c}\eta_{c}$ and $J/\psi J/\psi$ thresholds and is allowed to decay into both of these two final states. However, the width of the $J/\psi J/\psi$ final state is smaller, indicating that this state is more likely to decay into the $\eta_{c}\eta_{c}$ channel.
For the $1^{+-}$ and $2^{++}$ states, the masses are 6204.49 MeV and 6260.89 MeV, respectively. Both lie above the $\eta_{c}J/\psi$ and $J/\psi J/\psi$ thresholds. Nonetheless, due to quantum number constraints, the $1^{+-}$ state can only decay into the $\eta_{c}J/\psi$ channel, whereas the $2^{++}$ state is allowed to decay solely into the $J/\psi J/\psi$ channel.
The $P$-wave fully charmed tetraquark states have quantum numbers $J^{PC}=0^{-+}$, $1^{--}$, and $1^{-+}$. There is one $0^{-+}$ state, with a mass of 6543.23 MeV, which can decay only into the $\eta_{c}\chi_{c0}$ channel. Three states are identified with $J^{PC}=1^{--}$, having masses of 6524.51 MeV, 6555.97 MeV, and 6740.27 MeV, respectively. The two lower-mass states, at 6524.51 MeV and 6555.97 MeV, lie above the $J/\psi\chi_{c0}$ threshold but below the $J/\psi\chi_{c1}$ threshold. Consequently, due to conservation laws, they are only allowed to decay into the $\eta_{c}h_{c}$ and $J/\psi\chi_{c0}$ channels. The masses of both two states are in good agreement with the mass of the $X(6600)$ observed by the CMS Collaboration, reported as $M=6638^{+43}_{-38}(\text{stat})^{+16}_{-31}(\text{syst})$ MeV [7]. The highest $1^{--}$ state, at 6740.27 MeV, lies above the $J/\psi\chi_{c2}$ threshold and can decay into $\eta_{c}h_{c}$, $J/\psi\chi_{c0}$, $J/\psi\chi_{c1}$, and $J/\psi\chi_{c2}$ final states. This state is more likely to decay into the $\eta_{c}h_{c}$ and $J/\psi\chi_{c0}$ channels due to their larger widths. In addition, there is a single $1^{-+}$ state with a mass of 6586.89 MeV, which can decay only into the $\eta_{c}\chi_{c1}$ channel.
@@ III.2.2 $bb][\bar{b}\bar{b}]$ system
The quantum numbers of the $S$-wave fully bottom tetraquark states are $J^{PC}=0^{++}$, $1^{+-}$, and $2^{++}$. There are two $0^{++}$ states with masses of 18948.9 MeV and 19066.9 MeV. The $1^{+-}$ and $2^{++}$ states have masses of 19006.9 MeV and 19037.7 MeV, respectively. All these $S$-wave states lie above the $\Upsilon\Upsilon$ threshold. Both $0^{++}$ states are allowed to decay into the $\eta_{b}\eta_{b}$ and $\Upsilon\Upsilon$ channels.But this state is more likely to decay into the $\eta_{b}\eta_{b}$ channel due to its larger width. The $1^{+-}$ state can decay only into the $\eta_{b}\Upsilon$ channel, while the $2^{++}$ state is allowed to decay only into the $\Upsilon\Upsilon$ channel.
The quantum numbers of the $P$-wave fully bottom tetraquark states are $J^{PC}=0^{-+}$, $1^{--}$, and $1^{-+}$. Their masses all lie above the $\Upsilon\chi_{b2}$ threshold. Taking conservation laws into account, the $0^{-+}$ state can decay into the $\eta_{b}\chi_{b0}$ and $\Upsilon h_{b}$ channels, while the $1^{-+}$ state is allowed to decay into the $\eta_{b}\chi_{b1}$ and $\Upsilon h_{b}$ channels. The $1^{--}$ state can decay into multiple channels, including $\eta_{b}h_{b}$, $\Upsilon\chi_{b0}$, $\Upsilon\chi_{b1}$, and $\Upsilon\chi_{b2}$.
@@ III.2.3 $cb][\bar{c}\bar{b}]$ system
Since the $cb][\bar{c}\bar{b}]$ tetraquark system contains no identical quarks or antiquarks, it is not subject to the Pauli exclusion principle. As a result, there is no restriction on the construction of basis states, and all possible spin-parity and charge-conjugation quantum numbers can be realized. This leads to a rich spectra of possible states, including $0^{++}$, $^{++}, 1+1^{+-}$, $2^{++}$ for $S$-wave and $^{--}, 0+0^{-+}$, $^{--}, 1+1^{-+}$ for $P$-wave.
@@
There are four states with quantum numbers $0^{++}$, with masses of 12477.8 MeV, 12548.6 MeV, 12633.4 MeV, and 12690.1 MeV, respectively. Among them, the two lighter states, 12477.8 MeV and 12548.6 MeV, lie above the $\eta_{c}\eta_{b}$ threshold but below the $J/\psi\Upsilon$ threshold. Therefore, considering the conservation of angular momentum and $C$-parity, they can only decay into the $\eta_{c}\eta_{b}$ channel. In addition, the $cb][\bar{c}\bar{b}]$ tetraquark states can decay not only into $c\bar{c}$ and $b\bar{b}$ mesons, but also into $c\bar{b}$ and $b\bar{c}$ mesons. However, since the masses of the two lighter states are below the $B_{c}\bar{B}_{c}$ threshold, decays into $c\bar{b}$ and $b\bar{c}$ mesons are kinematically forbidden. The two heavier states, with masses of 12633.4 MeV and 12690.1 MeV, lie above the $J/\psi\Upsilon$ threshold and can thus decay into both $\eta_{c}\eta_{b}$ and $J/\psi\Upsilon$ channels. Moreover, since their masses are also above the $B_{c}\bar{B}_{c}$ threshold, decays into this channel are allowed. In particular, the state with mass 12690.1 MeV lies above the $B_{c}^{*}\bar{B}_{c}^{*}$ threshold, and therefore can also decay into the $B_{c}^{*}\bar{B}_{c}^{*}$ channel. The two states, 12633.4 MeV and 12690.1 MeV, are more likely to decay into $\eta_{c}\eta_{b}$ channel due to its larger width.
@@
There are two states with quantum numbers $1^{++}$, with masses of 12582.8 MeV and 12665.6 MeV, respectively. Both states lie above the $J/\psi\Upsilon$ threshold and are therefore allowed to decay into this channel. In addition, the heavier state, with a mass of 12665.6 MeV, lies above the thresholds for both $B_{c}\bar{B}_{c}^{*}$ and $B_{c}^{*}\bar{B}_{c}^{*}$, making decays into these channels kinematically allowed as well. But the width of $J/\psi\Upsilon$ channel is larger than $B_{c}\bar{B}_{c}^{*}$ and $B_{c}^{*}\bar{B}_{c}^{*}$, thus the state of 12665.6 MeV is more likely to decay into $J/\psi\Upsilon$ channel.
@@
There are four states with quantum numbers $1^{+-}$, with masses of 12547.4 MeV, 12590.8 MeV, 12617.6 MeV, and 12655.4 MeV, respectively. All of these states lie above the $\eta_{c}\Upsilon$ and $J/\psi\eta_{b}$ thresholds, allowing them to decay into both channels. Additionally, the two heavier states, with masses of 12617.6 MeV and 12655.4 MeV, are above the $B_{c}\bar{B}_{c}^{*}$ threshold and can therefore also decay into this channel. Both the four $1^{+-}$ states are more likely to decay into $\eta_{c}\Upsilon$ channel because the width of $\eta_{c}\Upsilon$ is larger than that of other channels.
@@
There are two states with quantum numbers $2^{++}$, whose masses are 12649.1 and 12666.1 MeV, respectively. Both them lie above the $J/\psi\Upsilon$ threshold and are allowed to decay into this channel. Furthermore, the state with 12666.1 MeV lies above the threshold of $B_{c}^{*}\bar{B}_{c}^{*}$, therefore can also decay into the $B_{c}^{*}\bar{B}_{c}^{*}$ channel. The state of 12666.1 MeV is more likely to decay into $J/\psi\Upsilon$ channel due to the larger width.
@@
There are two $0^{--}$ states, whose masses are 12955.1 MeV and 13037.9 MeV. The mass of 13037.9 MeV lies above $J/\psi\chi_{b1}$, $\chi_{c1}\Upsilon$ and $B_{c}\bar{B}_{c0}$ thresholds therefore this state can naturally decay into these channels. The state with a mass of 12955.1 MeV has no available $S$-wave decay channels. Because its mass lies below $J/\psi\chi_{b1}$, $\chi_{c1}\Upsilon$ and $B_{c}\bar{B}_{c0}$ thresholds. However, it can decay into the $J/\psi\Upsilon$ channel via a $P$-wave transition.
@@
There are four $0^{-+}$ states with masses of 12919.7 MeV, 12963.1 MeV, 12989.9 MeV and 13027.7 MeV. Their masses lie all above the thresholds of $\eta_{c}\chi_{b0}$ and $\chi_{c0}\eta_{b}$, thus they can naturally decay into these two channels. The heaviest state lies above the thresholds of the $J/\psi h_{b}$, $h_{c}\Upsilon$, and $B_{c}\bar{B}_{c0}$ channels, and can therefore decay spontaneously into all three of them. Both the four states are more likely to decay into $\chi_{c0}\eta_{b}$ channel due to the larger width.

### v2

@@ III.1 Model parameters
$M=$ $2m_{Q_{1}}+m_{Q_{2}}+\left(\kappa_{Q_{1}Q_{1}}\right)^{\bar{3}}\left[\bm{S}\left(\bm{S}+1\right)-\frac{3}{2}\right]$
$+$ $\left(\kappa_{Q_{1}Q_{2}}\right)^{\bar{3}}\left[J\left(J+1\right)-\bm{S}\left(\bm{S}+1\right)-\frac{3}{4}\right],$ (14)
where $J$ is the total angular momentum quantum number of the baryon, and $\bm{S}$ represents the total spin of the two identical-flavor quarks within the baryon. Therefore, we can obtain the model parameters $\kappa^{\bar{3}}$ required for the calculation. All the parameterss are shown in Table 1.
For the tetraquark configuration $|(Q_{1}Q_{2})(\bar{Q}_{3}\bar{Q}_{4})\rangle$, its diquark-antidiquark color characteristic is a superposition of $|(Q_{1}Q_{2})^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})^{3}\rangle$ and $|(Q_{1}Q_{2})^{6}(\bar{Q}_{3}\bar{Q}_{4})^{\bar{6}}\rangle$. According to the hypothesis of strict Casimir scaling, in the configuration $|(Q_{1}Q_{2})^{\bar{3}}(\bar{Q}_{3}\bar{Q}_{4})^{3}\rangle$, the parameter $(\kappa_{Q\bar{Q}})^{8}=\frac{1}{4}(\kappa_{Q\bar{Q}})^{1}$. Similarly, in the configuration $|(Q_{1}Q_{2})^{6}(\bar{Q}_{3}\bar{Q}_{4})^{\bar{6}}\rangle$, the parameters $(\kappa_{QQ})^{6}=-\frac{1}{2}(\kappa_{QQ})^{\bar{3}}$ and $(\kappa_{Q\bar{Q}})^{8}=\frac{5}{8}(\kappa_{Q\bar{Q}})^{1}$.
After studying the ground state of the fully heavy tetraquark state, we can include the contributions from $H_{SL}$ and $H_{L}$ to calculate the masses of the $L=1$ excited states. By applying Eq. (1) to charmonium mesons with $S=1$ and $L=0,1$, specifically $J/\psi$, $\chi_{c0}$, $\chi_{c1}$, and $\chi_{c2}$, we can fit the coefficients in $H_{SL}$ and $H_{L}$. We get $A_{cc}=21.83$ MeV and $B_{cc}=426.06$ MeV. Similarly, using the bottomonium mesons $\Upsilon$, $\chi_{b0}$, $\chi_{b1}$, and $\chi_{b2}$, we can obtain $A_{bb}=8.23$ MeV and $B_{bb}=438.72$ MeV. For the mixed charm-bottom tetraquark states, we use the average values of $A_{cc}$ and $A_{bb}$ (as well as $B_{cc}$ and $B_{bb}$) for our calculations.
Since the tetraquark states discussed in this work are structured as diquark–antidiquark systems, quark rearrangement is required for them to strongly decay into two mesons. The decay widths are calculated using Eq. (11) with $\mathcal{A}$ = 2600 MeV. The parameter $\mathcal{A}$ can be obtained by fitting the decay process of the scalar meson [Maiani:2004uc]. As higher-order decays (such as $P$-wave and $D$-wave decays) contribute less significantly to the total decay width, our analysis focuses primarily on $S$-wave decays. Of course, we also analyze and discuss $P$-wave decay modes in special cases.
@@
Figure 1: The mass spectra of the fully heavy tetraquark states: (a) $cc][\bar{c}\bar{c}]$, (b) $bb][\bar{b}\bar{b}]$, and (c) $cb][\bar{c}\bar{b}]$, each with different quantum numbers. The dash-dotted lines represent the corresponding meson-meson thresholds. In Fig. (c), the “several channels” include many threshold lines; for clarity and visual neatness, those lines overlapping with the energy levels have been omitted.
@@
Table 2: The quantum numbers, masses, eigenvectors of the basis wave functions, and decay channels of the fully heavy tetraquark $cc][\bar{c}\bar{c}]$ and $bb][\bar{b}\bar{b}]$ systems.
@@
Table 3: The quantum numbers, masses, eigenvectors of the basis wave functions, and decay channels of the fully heavy tetraquark $cb][\bar{c}\bar{b}]$ system. For some tetraquark states, there is no available decay channel that both satisfies the conservation law and has a mass higher than the corresponding meson-meson threshold.
@@
Table 4: The quantum numbers, masses, eigenvectors of the basis wave functions, and decay channels of $cc][\bar{c}\bar{b}]$ system.
@@
Table 5: The quantum numbers, masses, eigenvectors of the basis wave functions, and decay channels of $bb][\bar{b}\bar{c}]$ system.
@@
Table 6: The quantum numbers, masses, eigenvectors of the basis wave functions, and decay channels of the fully heavy tetraquark $cc][\bar{b}\bar{b}]$ system. Similarly, for certain states, there is no available decay channel that both satisfies the conservation law and has a mass higher than the corresponding meson-meson threshold.
@@
III.2 $cc][\bar{c}\bar{c}]$, $bb][\bar{b}\bar{b}]$, and $cb][\bar{c}\bar{b}]$ systems
From the basis wave functions of tetraquark $cc][\bar{c}\bar{c}]$, $bb][\bar{b}\bar{b}]$, and $cb][\bar{c}\bar{b}]$ configurations, it can be seen that these states possesses a well-defined charge conjugation parity ($C$-parity). For these states, conservation of $C$-parity must be taken into account during the decay process. Their mass spectra and decay channels were shown in Fig. 1 and Tables 2-3. The two final-state mesons we choose for these decays are in the $1S$ and $1P$ states. To date, only the spin-0 states of the $B_{c}(1S)$ and $B_{c}(2S)$ mesons have been observed experimentally. Therefore, for the remaining $c\bar{b}$ mesons in the $1S$ and $1P$ states, we use the masses predicted by lattice QCD calculations [PhysRevLett.121.202002]. For simplicity, we denote the $c\bar{b}$ mesons with quantum numbers $J^{P}=0^{+}$ and $1^{+}$ as $B_{c0}$ and $B_{c1}$, respectively.
@@ III.2.1 $cc][\bar{c}\bar{c}]$ system
The $1S$-wave fully charmed tetraquark states possess quantum numbers $J^{PC}=0^{++}$, $1^{+-}$, and $2^{++}$. There are two states with quantum numbers $0^{++}$, and their masses are 6098.45 MeV and 6314.21 MeV, respectively. The lower-mass state, at 6098.45 MeV, lies above both the $\eta_{c}\eta_{c}$ and $\eta_{c}J/\psi$ thresholds. However, due to conservation laws, it can only decay naturally into the $\eta_{c}\eta_{c}$ channel, while decay into $\eta_{c}J/\psi$ is forbidden. The higher-mass state, at 6314.21 MeV, lies above both $\eta_{c}\eta_{c}$ and $J/\psi J/\psi$ thresholds and is allowed to decay into both of these two final states. However, the narrower width of the $J/\psi J/\psi$ final state suggests that this state is more likely to be found in the $J/\psi J/\psi$ invariant mass spectrum. Furthermore, the wide width of the decay to $\eta_{c}\eta_{c}$ indicates that this state is difficult to distinguish from the background.
For the fully charmed tetraquark states with quantum numbers $1^{+-}$ and $2^{++}$, the mass spectra are 6204.49 MeV and 6260.89 MeV, respectively. Both lie above the $\eta_{c}J/\psi$ and $J/\psi J/\psi$ thresholds. Nevertheless, due to quantum number constraints, the $1^{+-}$ state can only decay into the $\eta_{c}J/\psi$ channel. The state with quantum numbers $2^{++}$ is allowed to decay solely into the final-state mesons $J/\psi J/\psi$. In addition, the narrow width of the decay to $J/\psi J/\psi$ indicates that the state can be found within the invariant mass spectrum of $J/\psi J/\psi$. This suggests that the fully charmed tetraquark state with quantum number $2^{++}$ at 6260.89 MeV may be $X(6200)$ [Song:2024ykq], which has been confirmed in recent experimental analyses. Unfortunately, we did not find reliable candidates for $X(6600)$, $X(6900)$, and $X(7200)$ in the $1S$-wave fully charmed tetraquark system.
The $P$-wave fully charmed tetraquark states have quantum numbers $J^{PC}=0^{-+}$, $1^{--}$, and $1^{-+}$. For the quantum numbers $0^{-+}$, there exists only one tetraquark state with a mass of 6543.23 MeV. Clearly, the feature of this state being above the $\eta_{c}\chi_{c0}$ threshold allows it to naturally decay into two mesons $\eta_{c}\chi_{c0}$. Three states are identified with $J^{PC}=1^{--}$, having masses of 6524.51 MeV, 6555.97 MeV, and 6740.27 MeV, respectively. The two lower-mass states, at 6524.51 MeV and 6555.97 MeV, lie above the $J/\psi\chi_{c0}$ threshold but below the $J/\psi\chi_{c1}$ threshold. Consequently, due to conservation laws, they are only allowed to decay into the $\eta_{c}h_{c}$ and $J/\psi\chi_{c0}$ channels. The highest $1^{--}$ state, at 6740.27 MeV, lies above the $J/\psi\chi_{c2}$ threshold and can decay into $\eta_{c}h_{c}$, $J/\psi\chi_{c0}$, $J/\psi\chi_{c1}$, and $J/\psi\chi_{c2}$ final states. Based on the properties of the final-state decay width, this state is more easily observed in the $J/\psi\chi_{c1}$ and $J/\psi\chi_{c2}$ channels compared to the $\eta_{c}h_{c}$ and $J/\psi\chi_{c0}$ channels. In addition, there is a single $1^{-+}$ state with a mass of 6586.89 MeV, which can decay only into the $\eta_{c}\chi_{c1}$ channel.
@@ III.2.2 $bb][\bar{b}\bar{b}]$ system
The quantum numbers of the $1S$-wave fully bottom tetraquark states are $J^{PC}=0^{++}$, $1^{+-}$, and $2^{++}$. There are two $0^{++}$ states with masses of 18948.9 MeV and 19066.9 MeV. The $1^{+-}$ and $2^{++}$ states have masses of 19006.9 MeV and 19037.7 MeV, respectively. All these $S$-wave states lie above the $\Upsilon\Upsilon$ threshold. Both $0^{++}$ states are allowed to decay into the $\eta_{b}\eta_{b}$ and $\Upsilon\Upsilon$ channels. But this state is more likely to decay into the $\eta_{b}\eta_{b}$ channel due to its larger width. Correspondingly, these two states are more easily found in the $\Upsilon\Upsilon$ invariant mass spectrum. The $1^{+-}$ state can decay only into the $\eta_{b}\Upsilon$ channel, while the $2^{++}$ state is allowed to decay only into the $\Upsilon\Upsilon$ channel.
The quantum numbers of the $P$-wave fully bottom tetraquark states are $J^{PC}=0^{-+}$, $1^{--}$, and $1^{-+}$. Their masses all lie above the $\Upsilon\chi_{b2}$ threshold. Taking conservation laws into account, the $0^{-+}$ state can decay into the $\eta_{b}\chi_{b0}$ and $\Upsilon h_{b}$ channels, while the $1^{-+}$ state is allowed to decay into the $\eta_{b}\chi_{b1}$ and $\Upsilon h_{b}$ channels. The $1^{--}$ state can decay into multiple channels, including $\eta_{b}h_{b}$, $\Upsilon\chi_{b0}$, $\Upsilon\chi_{b1}$, and $\Upsilon\chi_{b2}$ channels.
@@ III.2.3 $cb][\bar{c}\bar{b}]$ system
Since the tetraquark $cb][\bar{c}\bar{b}]$ system contains no identical quarks or antiquarks, it is not subject to the Pauli exclusion principle. As a result, there is no restriction on the construction of basis wave functions, and all possible spin-parity and charge-conjugation quantum numbers can be realized. This leads to a rich spectra of possible tetraquark states, including $0^{++}$, $^{++}, 1+1^{+-}$, $2^{++}$ for $S$-wave and $^{--}, 0+0^{-+}$, $^{--}, 1+1^{-+}$ for $P$-wave.
@@
There are four states with quantum numbers $0^{++}$, and their masses are 12477.8 MeV, 12548.6 MeV, 12633.4 MeV, and 12690.1 MeV, respectively. Among them, the two lighter states, 12477.8 MeV and 12548.6 MeV, lie above the $\eta_{c}\eta_{b}$ threshold but below the $J/\psi\Upsilon$ threshold. Therefore, considering the conservation of angular momentum and $C$-parity, they can only decay into the $\eta_{c}\eta_{b}$ channel. In addition, the $cb][\bar{c}\bar{b}]$ tetraquark states can decay not only into $c\bar{c}$ and $b\bar{b}$ mesons, but also into $c\bar{b}$ and $b\bar{c}$ mesons. However, since the masses of the two lighter states are below the $B_{c}\bar{B}_{c}$ threshold, decays into $c\bar{b}$ and $b\bar{c}$ mesons are kinematically forbidden. The two heavier states, with masses of 12633.4 MeV and 12690.1 MeV, lie above the $J/\psi\Upsilon$ threshold and can thus decay into both $\eta_{c}\eta_{b}$ and $J/\psi\Upsilon$ channels. Moreover, since their masses are also above the $B_{c}\bar{B}_{c}$ threshold, decays into this channel are allowed. In particular, the state with mass 12690.1 MeV lies above the $B_{c}^{*}\bar{B}_{c}^{*}$ threshold, and therefore can also decay into the $B_{c}^{*}\bar{B}_{c}^{*}$ channel. The two states are more likely to decay into mesons $\eta_{c}\eta_{b}$ due to their larger width. Similarly, due to the narrower decay width, 12633.4 MeV is more easily detected in the $J/\psi\Upsilon$ invariant mass spectrum, while 12690.1 MeV is more easily detected in the $B_{c}^{*}\bar{B}_{c}^{*}$ invariant mass spectrum.
@@
For the quantum numbers $1^{++}$, there are two states with masses of 12582.8 MeV and 12665.6 MeV, respectively. Both states lie above the $J/\psi\Upsilon$ threshold and are therefore allowed to decay into this channel. In addition, the heavier 12665.6 MeV lies above the thresholds for both $B_{c}\bar{B}_{c}^{*}$ and $B_{c}^{*}\bar{B}_{c}^{*}$, making decays into these channels kinematically allowed as well. But the width of $J/\psi\Upsilon$ channel is larger than $B_{c}\bar{B}_{c}^{*}$ and $B_{c}^{*}\bar{B}_{c}^{*}$, thus the state of 12665.6 MeV is more likely to decay into $J/\psi\Upsilon$ channel. Of course, the significantly narrower width indicates that this state is more likely to be observed in the $B_{c}^{*}\bar{B}_{c}^{*}$ invariant mass spectrum.
@@
For the quantum numbers $1^{+-}$, there are four states with masses of 12547.4 MeV, 12590.8 MeV, 12617.6 MeV, and 12655.4 MeV, respectively. All of these states lie above the $\eta_{c}\Upsilon$ and $J/\psi\eta_{b}$ thresholds, allowing them to decay into both channels. Based on decay width, 12582.8 MeV is more likely to be found in the $J/\psi\eta_{b}$ invariant mass spectrum, while 12665.6 MeV is more likely to be found in the $B_{c}\bar{B}_{c}^{*}$ invariant mass spectrum. Additionally, the two heavier states, with masses of 12617.6 MeV and 12655.4 MeV, are above the $B_{c}\bar{B}_{c}^{*}$ threshold and can therefore also decay into this channel. Both the four $1^{+-}$ states are more likely to decay into mesons $\eta_{c}\Upsilon$ because the width of $\eta_{c}\Upsilon$ is larger than that of other channels. Similarly, 12617.6 MeV is more readily observed in the $B_{c}\bar{B}_{c}^{*}$ invariant mass spectrum.
@@
There are two states with the quantum numbers $2^{++}$, whose masses are 12649.1 and 12666.1 MeV, respectively. Both them lie above the $J/\psi\Upsilon$ threshold and are allowed to decay into this channel. Furthermore, the state with 12666.1 MeV lies above the threshold of $B_{c}^{*}\bar{B}_{c}^{*}$, therefore can also decay into the $B_{c}^{*}\bar{B}_{c}^{*}$ channel. The 12666.1 MeV is more likely to decay into mesons $J/\psi\Upsilon$ due to the larger width. The smaller width makes this state easier to observe in the $B_{c}^{*}\bar{B}_{c}^{*}$ invariant mass spectrum.
@@
There are two $0^{--}$ states, whose masses are 12955.1 MeV and 13037.9 MeV. The mass of 13037.9 MeV lies above $J/\psi\chi_{b1}$, $\chi_{c1}\Upsilon$ and $B_{c}\bar{B}_{c0}$ thresholds therefore this state can naturally decay into these channels. Furthermore, due to its narrower width, this state is more easily detected in the $B_{c}\bar{B}_{c0}$ invariant mass spectrum. The state with a mass of 12955.1 MeV has no available $S$-wave strong decay channels. Because its mass lies below $J/\psi\chi_{b1}$, $\chi_{c1}\Upsilon$ and $B_{c}\bar{B}_{c0}$ thresholds. However, it can decay into the $J/\psi\Upsilon$ channel via a $P$-wave transition.
@@
For quantum numbers $0^{-+}$, there are four states with masses of 12919.7 MeV, 12963.1 MeV, 12989.9 MeV and 13027.7 MeV, respectively. Their masses lie all above the thresholds of $\eta_{c}\chi_{b0}$ and $\chi_{c0}\eta_{b}$, thus they can naturally decay into these two channels. The heaviest state lies above the thresholds of the $J/\psi h_{b}$, $h_{c}\Upsilon$, and $B_{c}\bar{B}_{c0}$ channels, and can therefore decay spontaneously into all three of them. The significantly narrower width suggests that 13027.7 MeV may form a peak in the $h_{c}\Upsilon$ invariant mass spectrum. Besides, both the four states are more likely to decay into mesons $\chi_{c0}\eta_{b}$ due to the larger width.